OpenStudy (anonymous):

write this quadratic equation in standard form. then use the quadratic formula to find the solutions. 12=2x squared -2x please show work :) need help

OpenStudy (anonymous):

2x

OpenStudy (accessdenied):

Standard form is ax^2 + bx + c = 0 so, just subtract 12 from both sides to get 2x^2 - 2x - 12 = 0 \[ Quadratic~Formula:~~x=\frac{-b \pm \sqrt{b^2 - 4ac}}{2a} \] a=2, b=-2, c=-12